ACTIVITY 7. EXPLORING MORE ELEMENTS OF HYPOTHESIS TESTING

A. Complete the table showing the rejection regions for common values of ๐œถ.

Alternative Hypothesis

๐œถ Left-Tailed Right-Tailed Two-Tailed ๐›ผ = .10 1. ___________ z > 1.28 z < -1.645 or z > 1.645

๐›ผ = .05 z < -1.645 2. _________ z < -1.96 or z > 1.96

๐›ผ = 0.25 3. ___________ z > 1.96 4. _______________

๐›ผ = .01 z < -2.33 5.__________ z < -2.58 or z > 2.58


1
Expert's answer
2022-05-09T17:54:39-0400

1.

The rejection region for right-tailed test for "\alpha=0.10" is

"R = \{z: z > 1.28\}"

The rejection region for two-tailed test for "\alpha=0.10" is

"R = \{z: z < -1.645\ or\ z>1.645\}"

2.

The rejection region for left-tailed test for "\alpha=0.05" is

"R = \{z: z <-1.645\}"

The rejection region for two-tailed test for "\alpha=0.05" is

"R = \{z: z < -1.96\ or\ z>1.96\}"

3.

The rejection region for right-tailed test for "\alpha=0.025" is

"R = \{z: z > 1.96\}"



4.

The rejection region for left-tailed test for "\alpha=0.01" is

"R = \{z: z <-2.33\}"



5. The rejection region for two-tailed test for "\alpha=0.01" is

"R = \{z: z < -2.58\ or\ z>2.58\}"